Q. 1. The demand function is q(p) = (p+1)-2  


a. What is the price elasticity of demand?

b. At what price is the price elasticity of demand equal to (-1)?

c. Write an expression for total revenue as a function of the price


1
Expert's answer
2022-01-13T09:12:05-0500

"a)"

Given that "g(p)=(p+1)^{-2}", the price elasticity of demand is given as,

"e_p={p\over g(p)}\times{g'(p)}"

"g'(p)=-2(p+1)^{-3}"

"e_p={p\over (p+1)^{-2}}\times -2(p+1)^{-3}=-{2p\over p+1}"

Therefore, the price elasticity of demand at price "p" is "-2p\over p+1"


"b)"

We set the price elasticity of demand found above equal to "-1" and solve for the price "p"

So,

"{-2p\over p+1}=-1\implies -2p=-p-1\implies p=1"

Therefore, when price=1, the price elasticity of demand is -1.


"c)"

The  total revenue is given as, "R(p)=g(p)\times p={1\over (p+1)^2}\times p={p\over (p+1)^2}". Therefore, the total revenue is given as,

"R(p)={p\over(p+1)^2}".